How many different 5-card poker hands would contain only cards of a single suit?
I know there are 52 cards in a deck and there are 4 different suits with 13 cards in each suit. From here, I am lost.

Answer by Alan3354(8169)

How many different 5-card poker hands would contain only cards of a single suit?
I know there are 52 cards in a deck and there are 4 different suits with 13 cards in each suit. From here, I am lost.
-----------------
It's 5 cards out of 13.
The 1st is 1 of 13, then 1 of 12, etc = 13*12*11*10*9
= 154,440
----------
But, since 2, 3, 4, 5, 6 is the same as 6,5,4,3,2 it's necessary to divide by 5*4*3*2*1 = 120
--> 154440/120 = 1287 possibilities.
-----------------
To be more formal, it's
13!/(5!*(13-5)!)
! is factorial. 5! = 5*4*3*2*1, etc.